A normal distribution has μ = 80 and σ = 10. What is the probability of randomly selecting between the mean and a score of 50


1
Expert's answer
2021-06-08T17:13:22-0400

Since  "\mu" =80 and  "\sigma=10" we have

We need to compute  Pr(X≤50). The corresponding z-value needed to be computed is:

"Z= \frac{(x-mean)}{sd}"


At mean , "Z= \frac{(80-80)}{10}=0"


At x= 50, "Z= \frac{(50-80)}{10}=-3"


P(z<0)=0.

P(z<-3.00)=0.00135


P(z<−3.00)=0.00135


P(50<x<80) =P(z<0)−P(z<−3.00)

=0.5−0.00135

=0.4986